SpaceRef

SpaceRef


NOAA SATOPS Morning Report: Tuesday, January 03, 2012

Status Report From: NOAA Office of Satellite Operations
Posted: Tuesday, January 3, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:

- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 003

GOES
GOES-12 DOY001 intermittent frame breaks from 03:09z through 07:07z

All other GOES operations were nominal over the past 96 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
N-17 DOY001 GACS 1A AND 2B REV. 49482 data lost due to an equipment problem at WCDA.

All other POES operations were nominal over the past 96 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
OPS47 DOY325 14:50Z through 21:20z the s/c was in a backup attitude mode configuration due to an ephemeris file upload problem. Under investigation.

All other DMSP operations were nominal over the past 96 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 96 hours.
No anomalies affecting product delivery or s/c health & safety

ESPC Reported data outage and delays
DOY001 N -17 GACS 1A and 2B REV. 49482 data lost due to an equipment problem at WCDA.

DOY001 ESPC did not receive N-17 GACS 1A and 2B on rev49482 due to an equipment problem at WCDA.

GOES-15 WEST 0100Z IMAGE LOST DUE TO GINI NPORT6 PRIMARY INGESTER HANGUP.

DOY002 at 1930z through DOY003 0300z, ESPC did not distribute the following Satellite images/products due to a ESPC Storage Area Network (SAN problem): Single Field of View ASOS Hourly, GOES LCB/ASH, GOES East and West Sounder, MET-9, GOES East and West Remapper, and HYDRO-Estimator).

On December 29, 2011 at 2300z until December 30, 2011 at 2040z, ESPC did not receive EOS MODIS/AIRS data due to network problem. Under investigation.

Questions or comments to Ron.Mahmot@noaa.gov

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.